# Learning to Think: Information-Theoretic Reinforcement Fine-Tuning for LLMs

### v2

@@ 5.2 Effectiveness and Efficiency Analysis
We compare L2T with the baselines across all the benchmarks and base models, recording both pass@1 accuracy and the required token budget. To reduce variance from limited samples, we use 20 outputs per question. Table 1 and Figure 2 show that L2T attains SOTA performance, achieving the highest reasoning effectiveness with the smallest token budget. For example, compared to outcome-reward-based methods, L2T delivers over a 3.7% gain in pass@1 and roughly doubles token efficiency; compared to methods focused on test-time compute, it achieves more than a 2% accuracy improvement while reducing the token budget by 20%. Moreover, L2T consistently outperforms baselines on multiple datasets with distributions different from the training data, further demonstrating its effectiveness across diverse tasks. These results confirm the superiority of our approach, which achieves effective reasoning with higher efficiency.
More efficient use of test-time compute Based on Subsection 5.1, we sample reasoning trajectories across various benchmarks with a fixed token context window. We truncate these trajectories at different token budgets and evaluate performance. Figure 3 shows the success rate against token consumption. We observe that (i) under the same token budget, L2T achieves higher reasoning accuracy; (ii) L2T consumes only 18% of the tokens required by the base model, 50% of those used by outcome-reward fine-tuning, and approximately 20% fewer tokens than process-reward models. These results demonstrate that L2T more effectively leverages test-time compute.
@@ 5.3 Ablation study
We conduct a series of ablation studies to evaluate the contribution of each component within L2T, the best parameterization and implementation choices, etc. See Appendix F for more results.
The effect of different components.
We evaluate three alternative configurations: (i) replacing the fitting information gain with a task-specific reward model; (ii) removing the parameter-compression penalty; and (iii) substituting the low-rank approximation with random sampling of 30% layers. Notably, the overall contribution of our reward has already been demonstrated in Subsection 5.2. From Figure 5 and Appendix F.3, we observe that both the fitting information gain and the parameter-compression penalty are critical for LLM reasoning; although random sampling is faster than low-rank approximation, it introduces additional error. These findings underscore the advantages of our design.
Parameter sensitivity.
We determine the hyperparameters of L2T by evaluating reasoning performance across benchmarks. Both $\alpha$ and $\beta$ are swept over $[0.3,0.9]$. We first use grid search to screen the parameters with a difference of 0.05, then refine it with 0.01, recording the average outcome. As shown in Figure 5, the optimal setting is $\alpha=0.8$ and $\beta=0.6$, also our choices.
@@ 6 Conclusion
  • Appendix A provides proofs of the theorem in the main text.
  • Appendix B provides the pseudo-code of our method.
  • Appendix C provides more discussions and further theoretical analysis about our work.
  • Appendix D provides the additional details of the benchmark datasets.
  • Appendix E provides the additional details of the implementation details.
  • Appendix F provides the full results and additional experiments.
Appendix A Proofs
In this section, we provide the proofs of Theorem 4.2 in the main text. The theoretical analyses of the proposed dense process reward are provided in Appendix C.1.
@@ Appendix A Proofs
$I(\tilde{\theta};s)=\mathbb{E}_{s}\Bigl{[}\mu_{KL}\bigl{(}p(\tilde{\theta}\mid s)\;\|\;p(\tilde{\theta})\bigr{)}\Bigr{]}.$ (6)
@@
Consider that the agents at step $k$ and step $(k-1)$ be centered on the same prior $\tilde{\theta}_{0}$, then we get:
@@
Proof.
For a given state $s_{k}$, the posterior distribution $p(\tilde{\theta}_{k}\mid s_{k})\propto p(s_{k}\mid\tilde{\theta}_{k})\;p(\tilde{\theta}_{k})$, take the logarithm to get the unstandardized posterior $L(\tilde{\theta}_{k})=\log p(s_{k}\mid\tilde{\theta}_{k})+\log p(\tilde{\theta}_{k})$. Among them,
@@
Let $\hat{\theta}=\arg\max_{\tilde{\theta}}L(\tilde{\theta})$ be the MAP (maximum a posteriori) estimate point of the posterior, and perform a second-order Taylor expansion of $L(\tilde{\theta})$ at $\hat{\theta}$ to obtain:
@@
Next, we discuss the relationship between Hessian and Fisher information. First, the observed Fisher of log-likelihood is defined as $-\sum_{i}\nabla^{2}_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}|s_{k})=\sum_{i}\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}|s_{k})\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}|s_{k})^{\top}$. When the third-order and higher derivatives are ignored, it can be approximated by:
$-H_{\rm lik}(\hat{\theta}_{k})\approx\sum_{i=1}^{n_{k}}\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}\mid s_{k})\,\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}\mid s_{k})^{\top}.$ (20)
@@
$-H(\hat{\theta}_{k})=-\bigl{(}H_{\rm lik}+H_{\rm prior}\bigr{)}\approx F_{\hat{\theta}_{k}},\quad\Sigma_{0}\approx(-H(\hat{\theta}_{k}))^{-1}\approx F_{\hat{\theta}_{k}}^{-1}.$ (22)
Thus, we complete the proof of the Lemma A.1.
@@
Appendix B Pseudo-Code
The pseudo-code of L2T is shown in Algorithm 1, mainly showing the steps on the basis of the implementation of base models. The main code is provided in the supplementary materials.
Algorithm 1 Pseudo-Code of L2T
[1] Input: Task distribution $P_{task}$; base model $\pi_{\theta}$;
Output: Fine-tuned $\pi_{\theta}$
for iteration = $1,\cdots,T$ do
     Sample tasks $X\sim P_{task}$
     for each $x\in X$ do
         for each episode $k$ do
              Sample $N$ rollouts $\{z_{k}^{i},y_{k}^{i}\}\}_{i=1}^{N}\sim\pi_{\theta}(\cdot|s_{k})$
              Compute average progress reward $r_{k}^{\rm prg}$ via Definition 4.1 and Theorem 4.2
              Compute average outcome reward $r_{k}^{\rm out}$ following [16]
         end for
     end for
     Update the policy $\pi_{\theta}$ via Eq.5
end for
return solution
Appendix C More Discussion
C.1 Theoretical Analyses about the Proposed Reward
In this subsection, we present the theoretical analysis for the proposed information-theoretic dense process reward. Specifically, we first explain why the increment in model prediction accuracy can be used to approximate $\Delta I_{k}$ (Proposition C.1). Next, we provide a theoretical analysis demonstrating that, under the current approximation, the computation method in Theorem 4.2 significantly reduces computational complexity (Theorem C.2), while the approximation error (Theorem C.3) remains bounded, thereby supporting the practical computation of the proposed reward.
@@ C.1 Theoretical Analyses about the Proposed Reward
In the context of LLMs, the fitting information gain quantifies the contribution of each optimization episode to the reasoning ability by tracking changes in the predicted correctness probability, i.e., the output distributions of $\pi_{\theta}$. In contrast, calculating the parameter-compression penalty is more complex: it involves estimating the mutual information increment between the model parameters $\theta$ and the historical context $s_{k}$. Direct computation of this increment is intractable in the large parameter space of LLMs. To overcome this challenge, we propose an efficient approximation in Theorem 4.2 to estimate the penalty term. Next, we demonstrate that this method significantly reduces the computational complexity (Theorem C.2) with limited approximation error (Theorem C.3) to support the computation of our proposed reward.
First, we prove that in Theorem 4.2, computing the parameter-compression penalty via low-rank approximation and Fisher matrix estimation achieves speedups of several orders of magnitude.
Theorem C.2.
Let the parameter dimension be $d$ and the rank cutoff value be $r$ ($r\ll d$). Compared to the original full non-approximate computation, estimating the parameter-compression penalty via Theorem 4.2 reduces the complexity of quadratic-form evaluations by a factor of $\Theta\bigl{(}(r/d)^{2}\bigr{)}$.
@@
Next, in the low-rank method, the original vector $\theta\in\mathbb{R}^{d}$ is first mapped to an $r$-dimensional subspace using truncated SVD (or randomized SVD). The main computation comes from the multiplication and addition of the $p\times r$ matrix, so the complexity is $\Theta(d,r^{2})$. Then, in this subspace, the gradient Jacobian $J=\nabla_{\tilde{\theta}}\log\pi_{\theta}$ is calculated and an approximate Fisher matrix $\tilde{F}=J^{\top}J$ is formed. Each entry requires $d$ multiplications and additions, and the overall complexity is still $\Theta(d,r^{2})$. Finally, evaluating the quadratic form $(\Delta\tilde{\theta})^{\top}\tilde{F},(\Delta\tilde{\theta})$ only requires multiplication and addition of the $r\times r$ matrix and the length $r$ vector, with a complexity of $\Theta(r^{2})$; if $\tilde{F}$ needs to be further decomposed, it will be $\Theta(r^{3})$. Therefore, the overall complexity of the approximate method can be expressed as
@@
where $d=\dim(\tilde{\theta})$ is the parameter dimension, the first term $\frac{M}{6}B^{3}$ comes from the third-order remainder of the second-order Taylor expansion, and the second term is derived from the matrix Hoeffding inequality or Bernstein inequality following Matrix-Concentration theory [46].
@@
Let $u=n\,t$, we get
[…]
@@
If the update amount is normalized (or assumed $B\leq 1$), it can be simplified to $\epsilon=\frac{M}{6}B^{3}+\sqrt{\frac{8\,\ln(2^{d}/\delta)}{N_{\tau}K}}$. Thus, we complete the proof of Theorem C.3.
@@
Therefore, we can conclude that the computation method in Theorem 4.2 significantly reduces computational complexity (Theorem C.2), while the approximation error (Theorem C.3) remains bounded, thereby supporting the practical computation of the proposed reward.
C.2 More Comparison

### v3

@@ 5.2 Effectiveness and Efficiency Analysis
We compare L2T with the baselines across all the benchmarks and base models, recording both pass@1 accuracy and the required token budget. To reduce variance from limited samples, we use 20 outputs per question. Table 1 and Figure 2 show that L2T attains SOTA performance, achieving the highest reasoning effectiveness with the smallest token budget. For example, compared to outcome-reward-based methods, L2T delivers over a 3.7% gain in pass@1 and roughly doubles token efficiency; compared to methods focused on test-time compute, it achieves more than a 2% accuracy improvement while reducing the token budget by 20%. Moreover, L2T consistently outperforms baselines on multiple datasets with distributions different from the training data, further demonstrating its effectiveness across diverse tasks. Besides, we also assess the performance of L2T on more base models with different scales, e.g., DeepSeek-R1-Distill-Qwen-7B and Qwen2-7B-Instruct, and more reasoning tasks, e.g., code generation tasks. Notably, mathematical reasoning and code generation serve as classic benchmarks for testing an LLM’s complex reasoning ability snell2024scalingllmtesttimecompute ; deepseekai2025deepseekr1incentivizingreasoningcapability ; lu2021codexglue . The results in Appendix G.2, demonstrate the advantage of L2T. These results confirm the superiority of our approach, which achieves effective reasoning with higher efficiency.
More efficient use of test-time compute Based on Subsection 5.1, we sample reasoning trajectories across various benchmarks with a fixed token context window. We truncate these trajectories at different token budgets and evaluate performance. Figure 3 shows the success rate against token consumption. We observe that (i) under the same token budget, L2T achieves higher reasoning accuracy; (ii) L2T consumes only 18% of the tokens required by the base model, 50% of those used by outcome-reward fine‑tuning, and approximately 20% fewer tokens than process-reward models. These results demonstrate that L2T more effectively leverages test‑time compute.
@@ 5.3 Ablation study
We conduct a series of ablation studies to evaluate the contribution of each component within L2T, the best parameterization and implementation choices, etc. See Appendix G for more results.
The effect of different components. We evaluate three alternative configurations: (i) replacing the fitting information gain with a task-specific reward model; (ii) removing the parameter-compression penalty; and (iii) substituting the low-rank approximation with random sampling of 30% layers. Notably, the overall contribution of our reward has already been demonstrated in Subsection 5.2. From Figure 5 and Appendix G.3, we observe that both the fitting information gain and the parameter-compression penalty are critical for LLM reasoning; although random sampling is faster than low-rank approximation, it introduces additional error. These findings underscore the advantages of our design.
Parameter sensitivity. We determine the hyperparameters of L2T by evaluating reasoning performance across benchmarks. Both $\alpha$ and $\beta$ are swept over $[0.3,0.9]$. We first use grid search to screen the parameters with a difference of 0.05, then refine it with 0.01, recording the average outcome. As shown in Figure 5, the optimal setting is $\alpha=0.8$ and $\beta=0.6$, also our choices.
@@ 6 Conclusion
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments. This work is supported by the National Natural Science Foundation of China (No. 62506355).
@@ Acknowledgements
  • Appendix A provides the list of notations.
  • Appendix B provides proofs of the theorem in the main text.
  • Appendix C provides the pseudo-code of our method.
  • Appendix D provides more discussions and further theoretical analysis about our work.
  • Appendix E provides the additional details of the benchmark datasets.
  • Appendix F provides the additional details of the implementation details.
  • Appendix G provides the full results and additional experiments.
Appendix A List of Notations
We list the definitions of all notations from the main text as follows:
  • $\square$
    Symbols of Problem Settings
    • $\bullet$
      $\mathcal{D}_{\rm train}=\{(x_{i},y^{*}_{i})\}_{i=1}^{N}$: training dataset.
    • $\bullet$
      $x_{i}$: $i$-th question.
    • $\bullet$
      $y^{*}_{i}$: the oracle reasoning trace that leads to the correct answer.
    • $\bullet$
      $\pi_{\theta}$: the LLM (treated as a stochastic policy).
    • $\bullet$
      $z_{0:T}\sim\pi_{\theta}(\cdot\mid x_{i})$: the output sequence for $x_{i}$.
    • $\bullet$
      $B_{\rm token}$: the token budget.
    • $\bullet$
      $r(x_{i},z_{0:T})$: the reward, which equals 1 if $z_{0:T}$ matches the oracle trace $y^{*}_{i}$.
  • $\square$
    Symbols of Problem Reformulation
    • $\bullet$
      $\mathcal{D}_{\rm train}=\{(x_{i},y^{*}_{i})\}_{i=1}^{N}$: training dataset.
    • $\bullet$
      $\pi_{\theta}$: the LLM (treated as a stochastic policy).
    • $\bullet$
      $z_{0:T}\sim\pi_{\theta}(\cdot\mid x_{i})$: the output sequence for $x_{i}$.
    • $\bullet$
      $B_{\rm token}$: the token budget.
    • $\bullet$
      $K$: the number of episodes.
    • $\bullet$
      $s_{k}=(x,z_{1:k-1})$: the state of $k$-th episode.
    • $\bullet$
      $z_{k}\sim\pi_{\theta}(\cdot\mid s_{k})$: the token sequence $z_{k}=(z_{k}^{1},\dots,z_{k}^{N_{k}})$ for the $k$-th episode of $x_{i}$.
    • $\bullet$
      $N_{k}$: the number of tokens in $z_{k}$.
    • $\bullet$
      $r^{\rm out}=r(x,z_{1:K})\in\{0,1\}$: the outcome reward indicating final correctness.
    • $\bullet$
      $r_{k}^{\rm prg}$: the dense process reward at episode $k$ (Eq.3).
    • $\bullet$
      $J_{r}(\cdot)$: the predicted correctness probability.
  • $\square$
    Information-Theoretic Quantities
    • $\bullet$
      $H(Y|X)$: conditional entropy of $Y$ given $X$ (uncertainty without parameters).
    • $\bullet$
      $H(Y|X,\theta)$: residual uncertainty of $Y$ given $X$ and parameters $\theta$.
    • $\bullet$
      $I(\theta;Y|X)$: conditional mutual information between $\theta$ and $Y$ given $X$.
    • $\bullet$
      $\Delta I_{k}$: fitting information gain contributed by episode $k$.
    • $\bullet$
      $I(\theta;s_{k})$: mutual information between parameters and context.
  • $\square$
    Parameters, Proxies, and Updates
    • $\bullet$
      $\theta,\,\theta_{k-1},\,\theta_{k}$: model parameters before/after episode $k$.
    • $\bullet$
      $\tilde{\theta},\,\tilde{\theta}_{k-1},\,\tilde{\theta}_{k}$: low-rank proxies of $\theta$ via SVD (retain top $r$ components).
    • $\bullet$
      $p(\tilde{\theta}_{k})=\mathcal{N}(\tilde{\theta}_{k}\mid\mu_{k},\Sigma_{k})$: Gaussian assumption for the low-rank proxy.
    • $\bullet$
      $r/d$: retained-rank ratio (e.g., $1\%\!\sim\!10\%$ for 1.5B; $0.1\%\!\sim\!1\%$ for 7B).
  • $\square$
    GRPO Integration and Token-Level Credit
    • $\bullet$
      $R_{i,k}=\tfrac{1}{K_{i}}r_{i}^{\text{out}}+\alpha\,r^{\text{prg}}_{i,k}$: per-episode reward for rollout $i$ (with process-weight $\alpha$).
    • $\bullet$
      $w_{i,t}\propto-\log p_{\theta_{\text{old}}}(z_{i,t}\mid s_{i,t})$: token weights by log-probability surprise.
    • $\bullet$
      $r_{i,t}$: token-level rewards obtained by distributing $R_{i,k}$ using $w_{i,t}$.
    • $\bullet$
      $\tilde{r}_{i}$: 95% truncated mean of token rewards for rollout $i$.
    • $\bullet$
      $\hat{A}_{i}=(\tilde{r}_{i}-\bar{\tilde{r}})/\sigma_{\tilde{r}}$: group-level advantage (standardized).
    • $\bullet$
      $A_{i,t}=\hat{A}_{i}\cdot\dfrac{r_{i,t}}{\tilde{r}_{i}}$: token-level advantage rescaled by relative contribution.
Appendix B Proofs
In this section, we provide the proofs of Theorem 4.2 in the main text. The theoretical analyses of the proposed dense process reward are provided in Appendix D.1.
@@ Appendix B Proofs
[…]
@@
Why a Gaussian distribution is a common and mild assumption under this context. During LLM training, parameter updates $\theta$ can be viewed as the accumulation of numerous small stochastic gradient steps. Each step introduces a small, random perturbation in parameter space, effectively acting as the sum of many independent random variables. When these perturbations are sufficiently numerous and diverse, the overall distribution of parameter changes tends toward a Gaussian distribution, as implied by the Central Limit Theorem. Therefore, when estimating the compression term, we model the low-rank surrogate $\tilde{\theta}$ (obtained via SVD) using a Gaussian distribution. This assumption aligns with covariance approximation techniques in the PAC-Bayes framework and Fisher information-based methods, allowing us to derive a closed-form expression for mutual information and significantly reduce computational cost.
@@
Consider that the agents at step $k$ and step (k1)(k-1) be centered on the same prior $\tilde{\theta}_{0}$, then we get:
@@
Proof. For a given state $s_{k}$, the posterior distribution $p(\tilde{\theta}_{k}\mid s_{k})\propto p(s_{k}\mid\tilde{\theta}_{k})\;p(\tilde{\theta}_{k})$, take the logarithm to get the unstandardized posterior $L(\tilde{\theta}_{k})=\log p(s_{k}\mid\tilde{\theta}_{k})+\log p(\tilde{\theta}_{k})$. Among them,
@@
[…]
@@
Next, we discuss the relationship between Hessian and Fisher information. First, the observed Fisher of log-likelihood is defined as $\sum_{i}\nabla^{2}_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}|s_{k})=\sum_{i}\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}|s_{k})\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}|s_{k})^{\top}$. When the third-order and higher derivatives are ignored, it can be approximated by:
$H_{\rm lik}(\hat{\theta}_{k})\approx\sum_{i=1}^{n_{k}}\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}\mid s_{k})\,\nabla_{\tilde{\theta}_{k}}\log\pi_{\theta}(z_{i}\mid s_{k})^{\top}.$ (20)
@@
$H(\hat{\theta}_{k})=-\bigl(H_{\rm lik}+H_{\rm prior}\bigr)\approx F_{\hat{\theta}_{k}},\quad\Sigma_{0}\approx(-H(\hat{\theta}_{k}))^{-1}\approx F_{\hat{\theta}_{k}}^{-1}.$ (22)
Thus, we complete the proof of the Lemma B.1.
@@
Appendix C Pseudo-Code
The pseudo-code of L2T is shown in Algorithm 1, providing the main steps of L2T with GRPO.
Algorithm 1 Pseudo-Code of L2T (GRPO Version)
0: Initial policy $\pi_{\theta}$; prompt distribution $\mathcal{D}$; hyperparameters $\alpha$ and $\beta$
1:for step =1=1 to NN do
2:  Sample a batch $D_{b}$ from $\mathcal{D}$
3:  Set old policy $\pi_{\theta_{\rm old}}\leftarrow\pi_{\theta}$
4:  for each query $x_{i}\in D_{b}$ do
5:   Sample NN rollouts $\{y_{0},y_{1},\cdots,y_{N-1}\}\sim\pi_{\theta_{\rm old}}(\cdot\mid x_{i})$
6:   for each rollout $y_{i}$ do
7:    Compute $r_{i,k}^{\rm prg}$ via Definition 4.1 and Theorem 4.2 for each episode $k$
8:    Compute $r^{\rm out}_{i}$ following [16]
9:    Compute per-episode reward $R_{i,k}=\tfrac{1}{K_{i}}r_{i}^{\text{out}}+\alpha\,r^{\text{prg}}_{i,k}$
10:    Set surprise weights $w_{i,t}\propto-\log p_{\theta_{\text{old}}}(z_{i,t}\mid s_{i,t})$
11:    Assign per-token rewards $r_{i,t}\leftarrow\frac{w_{i,t}}{\sum_{t^{\prime}\in k}w_{i,t^{\prime}}}\cdot R_{i,k}$
12:   end for
13:   Compute truncated mean $\tilde{r}_{i}\leftarrow\mathrm{TruncMean}_{95\%}\big(\{r_{i,t}\}_{t}\big)$
14:   Compute group-level advantage $\hat{A}_{i}=\frac{\tilde{r}_{i}-\bar{\tilde{r}}}{\sigma_{\tilde{r}}}$
15:   Rescale to token-level advantages $A_{i,t}=\hat{A}_{i}\cdot\frac{r_{i,t}}{\tilde{r}_{i}}$
16:  end for
17:  Update $\pi_{\theta}$ via the GRPO objective
18:end for
19:return $\pi_{\theta}$
Appendix D More Discussion
D.1 Theoretical Analyses about the Proposed Reward
In this subsection, we present the theoretical analysis for the proposed information-theoretic dense process reward. Specifically, we first explain why the increment in model prediction accuracy can be used to approximate $\Delta I_{k}$ (Proposition D.1). Next, we provide a theoretical analysis demonstrating that, under the current approximation, the computation method in Theorem 4.2 significantly reduces computational complexity (Theorem D.2), while the approximation error (Theorem D.3) remains bounded, thereby supporting the practical computation of the proposed reward.
@@ D.1 Theoretical Analyses about the Proposed Reward
In the context of LLMs, the fitting information gain quantifies the contribution of each optimization episode to the reasoning ability by tracking changes in the predicted correctness probability, i.e., the output distributions of $\pi_{\theta}$. In contrast, calculating the parameter-compression penalty is more complex: it involves estimating the mutual information increment between the model parameters $\theta$ and the historical context $s_{k}$. Direct computation of this increment is intractable in the large parameter space of LLMs. To overcome this challenge, we propose an efficient approximation in Theorem 4.2 to estimate the penalty term. Next, we demonstrate that this method significantly reduces the computational complexity (Theorem D.2) with limited approximation error (Theorem D.3) to support the computation of our proposed reward.
First, we prove that in Theorem 4.2, computing the parameter‑compression penalty via low‑rank approximation and Fisher matrix estimation achieves speedups of several orders of magnitude.
Theorem D.2
Let the parameter dimension be $d$ and the rank cutoff value be $r$ ($r\ll d$). Compared to the original full non‑approximate computation, estimating the parameter‑compression penalty via Theorem 4.2 reduces the complexity of quadratic‑form evaluations by a factor of $\Theta\bigl((r/d)^{2}\bigr)$.
@@
[…]
@@
[…]
@@
Let u=ntu=n\,t, we get
$\Pr\bigl(\|\widehat{F}-F\|\geq t\bigr)=\Pr\bigl(\|n(\widehat{F}-F)\|\geq n\,t\bigr)\;\leq\;2\,d\,\exp\!\Bigl(-\frac{n^{2}t^{2}}{8\,n\,R^{2}}\Bigr)=2\,d\,\exp\!\Bigl(-\frac{n\,t^{2}}{8\,R^{2}}\Bigr).$ (35)
@@
If the update amount is normalized (or assumed $B\leq 1$), it can be simplified to $\epsilon=\frac{M}{6}B^{3}+\sqrt{\frac{8\,\ln(2^{d}/\delta)}{N_{\tau}K}}$. Thus, we complete the proof of Theorem D.3.
@@
Therefore, we can conclude that the computation method in Theorem 4.2 significantly reduces computational complexity (Theorem D.2), while the approximation error (Theorem D.3) remains bounded, thereby supporting the practical computation of the proposed reward.
D.2 Intuition behind the Proposed Reward
D.2.1 Intuition behind the Fitting Information Gain
How to interpret LLM reasoning from an information-theoretic perspective (fitting gain vs. uncertainty)
The reasoning process of an LLM can be viewed as inferring the correct answer $Y$ from input $X$. The more certain the model’s prediction is, the better it “understands” the answer. Based on classical information theory [37, 3, 14], we can use conditional entropy $H(Y|X)$ to characterize the model’s uncertainty about the output $Y$. If the model is sufficiently confident, its $H(Y|X)$ should be low; conversely, if it is uncertain or making a blind guess, $H(Y|X)$ should be high. Within this framework, the goal of inference is to gradually reduce $H(Y|X)$ until the correct answer is output. Importantly, while $H(Y|X)$ represents the relationship between the input and output, this process is controlled by the LLM (determined by the parameter $\theta$). Therefore, a more reasonable metric is: Given $\theta$, what is the model’s uncertainty about $Y$? That is, we want $\theta$ to not only capture the input $X$ but also provide strong discrimination of the correct answer $Y$. Therefore, we use conditional mutual information $I(\theta;Y|X)=H(Y|X)-H(Y|X,\theta)$ to measure how much the known model parameters $\theta$ reduce the uncertainty about $Y$ given $X$. This ties rewards to meaningful reasoning progress, i.e., the reward for each episode depends on how much it helps the model understand the correct answer. Therefore, we define “fitting information gain as the reduction in uncertainty”.
How to characterize the gradual improvement of inference (introducing episode gain with $\theta_{k}$ and $\theta_{k-1}$)
In each episode $k$, the model updates its parameter state from $\theta_{k-1}$ to $\theta_{k}$ by observing the new reasoning step $z_{k}$. The key question is: Does this episode help the model better "understand" the answer? Therefore, we define fitting information gain of this episode as $\Delta I_{k}=I(\theta_{k};Y|X)-I(\theta_{k-1};Y|X)$. If this incremental gain is significant, it indicates that this episode has helped the model become more certain and effective. Here, $\theta_{k}$ and $\theta_{k-1}$ represent the posterior model parameters before and after learning the knowledge from episode $k$, estimated by the change in log-probability during the forward prediction process.
Why the proposed metric corresponds to reducing uncertainty (approximating information gain with $J_{r}$)
Because directly calculating mutual information is too complex, we introduce an approximate metric $J_{r}(\cdot)$ to represent the model’s predicted probability of the correct answer. This metric improves as the model’s “confidence” increases. Therefore, the fitting gain is approximated as $\Delta I_{k}\approx J_{r}(\pi_{\theta}(\cdot\mid s_{k},z_{k}))-J_{r}(\pi_{\theta}(\cdot\mid s_{k}))$, with theoretical guarantees in Appendix D.1. This difference measures whether episode $k$ improves the model’s ability to predict the correct answer, reflecting the reduction process, i.e., the increasing reliability of reasoning.
D.2.2 Intuition behind the Compression Penalty
Why introducing compression penalty
In LLM reasoning, if each episode’s information causes a significant change, while that episode may provide information gain, it may also capture unnecessary details within that episode, leading to overfitting or redundant computation. We want the model to only learn information that contributes to the answer within the episode. Therefore, inspired by the information bottleneck theory, we introduce a compression penalty based on the fitting term to further improve efficiency. It measures the “information overhead” incurred by each episode from an information-theoretic perspective.
How is it measured (intuition behind the design)
If $\theta_{k}$ differs significantly from $\theta_{k-1}$, but the model’s prediction performance (i.e., fitting gain) improves only slightly, this step may “absorb redundant information”. Therefore, we use the mutual information increment $I(\theta_{k};s_{k})-I(\theta_{k-1};s_{k-1})$ between them with the fitting term to measure whether the information in episode $k$ introduces unnecessary overhead.
Why this is called compression
The term compression comes from the idea that a model should retain only the minimal amount of information sufficient to perform accurate reasoning. In our setting, the mutual information $I(\theta;s_{k})$ quantifies how much the information of this episode is encoded into the model parameters $\theta$. A larger value implies that the model has to “store” more bits to fit that episode, akin to using more storage in a compressed file. By penalizing the increase $I(\theta_{k};s_{k})-I(\theta_{k-1};s_{k-1})$, we explicitly discourage storing redundant information, promoting a more compact (i.e., compressed) internal representation. This aligns with principles from MDL and PAC-Bayes, where generalization is favored when the hypothesis (here, $\theta$) is simple and concise.
D.3 More Comparison